# Semantic Mapping in Indoor Embodied AI – A Survey on Advances, Challenges, and Future Directions

### v2

@@ 6.3 Summary
In this section we summarize the key challenges and the future scope of evaluation (Tab. 7). Extrinsic evaluation metrics have been widely explored in the literature and continue to evolve with the emergence of increasingly complex tasks. However, evaluating the quality of maps (intrinsic) in semantic SLAM and embodied AI continues to be understudied and presents two key challenges. First, collecting accurate, fine-grained, open-vocabulary ground-truth annotations for large, real-world environments is inherently difficult. This limitation makes it particularly challenging to assess important properties of the built maps such as semantic accuracy, consistency, and completeness of the maps. Without reliable ground truth, these evaluations often remain coarse or dataset-specific. Second, there is currently no standardized suite of evaluation metrics that can be consistently applied across different types of map structures (topological, spatial grid, dense geometric, etc.) or across different encoding schemes (explicit and implicit features). Together, these challenges highlight a critical gap in the field and point toward the need for developing scalable, generalizable map evaluation frameworks, offering an important direction for future research.
@@ 7 Challenges
Standardized evaluation framework. A significant challenge in semantic map building lies in the absence of standardized evaluation frameworks that can consistently benchmark performance across diverse tasks, environments and map representations, as we discuss at length in Sec. 6.2. While the community has focused more on task-specific extrinsic evaluation, limited progress have been made towards developing intrinsic measures to evaluate the map quality. This lack of standardized intrinsic evaluation hinders progress, as it becomes unclear whether improvements in downstream task success stem from better semantic understanding or other factors like better control policies. Developing general, interpretable and task-agnostic evaluation metrics is crucial, not only to track progress but also to facilitate cross-domain generalization. As open-vocabulary multi-modal queryable general-purpose semantic maps become increasingly central to reasoning in embodied AI and real-world robots, creating a unified benchmark suite that reflects both structural and semantic fidelity is essential for advancing the field.
@@ 8.5 Devising evaluation metrics
As we discuss in Sec. 6, the evaluating semantic maps in embodied AI research has received limited attention compared to assessing agent performance in downstream tasks. However, we believe that advancing the field requires a stronger emphasis on map evaluation using metrics such as accuracy, completeness, consistency, and robustness. Regardless of the downstream task, maps should be assessed on how well they capture semantic information (accuracy), their geometric and semantic coverage (completeness), their spatial and semantic reliability in dynamic environments (consistency), and their confidence and ability to handle uncertainty and noise (robustness). Establishing standardized evaluation metrics and frameworks for semantic maps remains a critical challenge with substantial opportunities for future research.
@@ 8.6 Summary
This section outlines promising future directions, as derived from the current challenges in semantic mapping in Sec. 7. As the field moves toward creating flexible, general-purpose, and queryable maps, researchers face open problems in balancing semantic richness with efficiency, ensuring scalability and robustness in dynamic environments, and developing better evaluation metrics that go beyond task performance to assess map quality directly. These provide a multitude of future research directions, which will help advance the field.
@@ Acknowledgements
The authors were supported by Canada CIFAR AI Chair grant and NSERC Discovery Grant. We also thank Manolis Savva, Yasutaka Furukawa, Tommaso Campari, Austin T. Wang, Xingguang Yan, Bernadette Bucher, Duy Ta and Sachini Herath for their valuable feedback on our paper.

### v3

@@ 6.3 Summary
In this section we summarize the key challenges and the future scope of evaluation (Tab.˜7). Extrinsic evaluation metrics have been widely explored in the literature and continue to evolve with the emergence of increasingly complex tasks. However, evaluating the quality of maps (intrinsic) in semantic SLAM and embodied AI continues to be understudied and presents two key challenges. First, collecting accurate, fine-grained, open-vocabulary ground-truth annotations for large, real-world environments is inherently difficult. This limitation makes it particularly challenging to assess important properties of the built maps such as semantic accuracy, consistency, and completeness of the maps. Without reliable ground truth, these evaluations often remain coarse or dataset-specific. Second, there is currently no standardized suite of evaluation metrics that can be consistently applied across different types of map structures (topological, spatial grid, dense geometric, etc.) or across different encoding schemes (explicit and implicit features). Together, these challenges highlight a critical gap in the field and point toward the need for developing scalable, generalizable map evaluation frameworks, offering an important direction for future research.
@@ 7 Challenges
Standardized evaluation framework. A significant challenge in semantic map building lies in the absence of standardized evaluation frameworks that can consistently benchmark performance across diverse tasks, environments and map representations, as we discuss at length in Sec.˜6.2. While the community has focused more on task-specific extrinsic evaluation, limited progress have been made towards developing intrinsic measures to evaluate the map quality. This lack of standardized intrinsic evaluation hinders progress, as it becomes unclear whether improvements in downstream task success stem from better semantic understanding or other factors like better control policies. Developing general, interpretable and task-agnostic evaluation metrics is crucial, not only to track progress but also to facilitate cross-domain generalization. As open-vocabulary multi-modal queryable general-purpose semantic maps become increasingly central to reasoning in embodied AI and real-world robots, creating a unified benchmark suite that reflects both structural and semantic fidelity is essential for advancing the field.
@@ 8.5 Devising evaluation metrics
As we discuss in Sec.˜6, the evaluating semantic maps in embodied AI research has received limited attention compared to assessing agent performance in downstream tasks. However, we believe that advancing the field requires a stronger emphasis on map evaluation using metrics such as accuracy, completeness, consistency, and robustness. Regardless of the downstream task, maps should be assessed on how well they capture semantic information (accuracy), their geometric and semantic coverage (completeness), their spatial and semantic reliability in dynamic environments (consistency), and their confidence and ability to handle uncertainty and noise (robustness). Establishing standardized evaluation metrics and frameworks for semantic maps remains a critical challenge with substantial opportunities for future research.
@@ 8.6 Summary
This section outlines promising future directions, as derived from the current challenges in semantic mapping in Sec.˜7. As the field moves toward creating flexible, general-purpose, and queryable maps, researchers face open problems in balancing semantic richness with efficiency, ensuring scalability and robustness in dynamic environments, and developing better evaluation metrics that go beyond task performance to assess map quality directly. These provide a multitude of future research directions, which will help advance the field.
@@ Acknowledgements
The authors were supported by a Canada CIFAR AI Chair grant and an NSERC Discovery Grant. We also thank Manolis Savva, Yasutaka Furukawa, Tommaso Campari, Austin T. Wang, Xingguang Yan, Bernadette Bucher, Duy Ta and Sachini Herath, and the anonymous reviewers for their valuable feedback on our paper.